OpenStudy (anonymous):

Find all the zeroes of the equation 3x^2-4=-x^4

OpenStudy (freckles):

so you have x^4+3x^2-4=0

OpenStudy (freckles):

this is a quadratic in terms of x^2

OpenStudy (freckles):

that is if you let x^2=u then you see we have a quadratic to solve

OpenStudy (freckles):

\[u^2+3u-4=0\] solve this as a first step